Calculate the enthalpy of the following reaction: (3 marks) 2Mg + O2 → 2MgO Given: Mg + 2HCl → MgCl2 + H2 ΔH = -21.0 kJ MgO + 2HCl → MgCl2 + H2 O ΔH = -33.5 kJ H2 + ½ O2 → H2 O ΔH = -76.0 kJ
1
Expert's answer
2020-06-04T10:31:59-0400

Solution:

Given:

(1) Mg + 2HCl → MgCl2 + H2, ΔH1 = -21.0 kJ

(2) MgO + 2HCl → MgCl2 + H2O, ΔH2 = -33.5 kJ

(3) H2 + ½O2 → H2O, ΔH3 = -76.0 kJ


2Mg + O2 → 2MgO ΔHr = ???


Using Hess's law to calculate the reaction enthalpy:

ΔHr = 2×ΔH1 - 2×ΔH2 + 2×ΔH3

ΔHr = 2×(-21.0 kJ) - 2×(-33.5 kJ) + 2×(-76.0 kL) = -127 kJ

ΔHr = -127 kJ


Answer: The enthalpy (ΔHr) of this reaction is -127 kJ.


NB! The values of reaction enthalpies (1), (2) and (3) are not correct!

(1) Mg + 2HCl → MgCl2 + H2, ΔH1 = -427.99 kJ/mol

(2) MgO + 2HCl → MgCl2 + H2O, ΔH2 = -74.66 kJ/mol

(3) H2 + ½O2 → H2O, ΔH3 = -285.5 kJ/mol

Then, ΔHr = -1277.66 kJ/mol